# Pushing the frontiers of subexponential FPT time for Feedback Vertex Set

### v1

@@ 3.3. Kernelization rules
Here we provide the four kernelization rules (KR1), (KR2), (KR3), and (KR4) that Algorithm 1 tries to apply Line 2. Each of these rules takes as input an instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS and outputs a single instance $(G^{\prime},k^{\prime},M^{\prime},\mathcal{H}^{\prime})$. Such a rule is said to be safe if:
  • On input an instance of $(r,\mathcal{G})$-Ann-FVS it returns an instance of $(r,\mathcal{G})$-Ann-FVS (in particular the graph of the output instance is a $K_{r,r}$-free graph in $\mathcal{G}$), and
@@
Given any instance $(G,k,M,\mathcal{H})$, recall that every connected component of $G-M$ is a tree. We root each of them at an arbitrary vertex. We define a subforest of $G-M$ as a subset $T\subseteq V(G)\setminus M$ and say the set is a subtree of $G-M$ if $G[T]$ is a tree. Given a vertex $v$ of a connected component $T$ in $G-M$, we define the subtree $T_{v}$ of $G-M$ as the connected component of $v$ in $G-M-u$, where $u$ is the parent of $v$, if any. If $v$ is the root of $T$, then $T_{v}=T$. In any case $T_{v}$ is rooted at $v$. Given $X\subseteq V(G)$, we denote $\mathcal{H}-X=\{H\in\mathcal{H}\mid H\cap X=\emptyset\}$. Given a subtree $T$ of $G-M$, $\partial_{\overline{M}}(T)$ denotes the set $\{v\in T,~{}N(v)\not\subseteq M\cup T\}$, and $b_{\overline{M}}(T)$ denotes the size of this set. We also denote $\operatorname{\mu}(T)=\max(d_{M}(T),b_{\overline{M}}(T))$.
We start with two basic reduction rules often used to deal with FVS and that allow to get rid of vertices of degree 1 and arbitrarily long paths of vertices of degree 2. Notice that we only apply here the reduction to vertices in $G-M$.
@@
    Given an instance $(G,k,M,\mathcal{H})$, if there exists a vertex $v\in V(G)\setminus M$ of degree $d(v)\leq 1$, output $(G-\{v\},k,M,\mathcal{H})$.
@@
    Given an instance $(G,k,M,\mathcal{H})$, if there exists a path $quvw$ in $V(G)\setminus M$ such that the four vertices have degree $2$ in $G$ and $d_{M}(u)=d_{M}(v)=0$, output $(G^{\prime},k,M,\mathcal{H})$, where $G^{\prime}$ is the graph obtained from $G$ by contracting the edge $uv$.
@@
Notice that the vertices considered in the two rules belong to $G-M$, in particular they do not belong to any member of $\mathcal{H}$ (which are subsets of $M$). Contracting an edge with endpoints of degree two preserves being $K_{r,r}$-free and does not modify the size of the minimum feedback vertex sets. Moreover the obtained graph is still in $\mathcal{G}$ by definition of a nice class. The output is then an instance of $(r,\mathcal{G})$-Ann-FVS equivalent to the input one. The running time claim is immediate.
@@
Observe that we did not use the condition in (KR2) that the endpoints of the considered path have degree at most $2$ in $G$. However it will later be used in 3.34 to bound the size of $M$ during the execution of the algorithm.
The next rule ensures that the vertices outside $M$ have a small neighborhood in $M$. It increases the size of $M$, but in a controlled manner as we will see later in 3.34.
@@
    Given an instance $(G,k,M,\mathcal{H})$, if there is a vertex $v\in V(G)\setminus M$ such that $d_{M}(v)\geq t=2d_{r}$, with $d_{r}$ the value defined in 1.5, output $(G,k,M\cup\{v\},\mathcal{H})$.
@@
The fourth kernelization rule will ensure that the number of neighbors outside $M$ of a vertex $v\in V(G-M)$ is strongly correlated with the size of the neighborhood of its descendants in $M$. Consider an instance $(G,k,M,\mathcal{H})$ where none of the previous rules applies. Given a family $\mathcal{T}$ of disjoint subtrees of $G-M$, a tree $T\in\mathcal{T}$ is redundant (for $\mathcal{T}$) if for all $v\in M$ such that $d_{T}(v)\geq 2$, there exists $T^{\prime}\in\mathcal{T}$ with $T^{\prime}\neq T$ such that $d_{T^{\prime}}(v)\geq 2$.
@@
Consider a set $X\subseteq M$ and a set $\mathcal{T}$ of subtrees of $G-M$ with $|\mathcal{T}|\geq|X|+1$, and such that $N_{M}(T)=X$ for every $T\in\mathcal{T}$. Then, there exists a redundant tree for $\mathcal{T}$, and it can be found in polynomial time.
@@
For any $x\in X$, if there is at least one subtree $T\in\mathcal{T}$ such that $d_{T}(x)\geq 2$, we arbitrarily pick one of them and call it $T_{x}$. Now, let $T^{\prime}$ be one of the remaining trees in $\mathcal{T}$. Such a tree exists as $|\mathcal{T}|\geq|X|+1$, and there are at most $|X|$ trees for the form $T_{x}$, for some $x\in X$. It is immediate from the definition that $T^{\prime}$ is redundant. ∎
Recall that we consider a fixed child-parent orientation in the forest $G-M$. In what follows, we say that a subset $F\subseteq V(G-M)$ is a downward-closed subtree (or subforest when the set is not necessarily connected) of $G-M$ when for any $v\in F$ and any children $u$ of $v$, $u\in F$. The fourth kernelization rule is as follow:
@@
    Given an instance $(G,k,M,\mathcal{H})$, a set $X\subseteq M$ with $|X|\geq 1$ and a set $\mathcal{T}$ of at least $|X|+2$ disjoint downward-closed subtrees of $G-M$ such that:
@@
    • either all the roots of the trees in $\mathcal{T}$ have a common parent $r$, or $\mathcal{T}$ consists only in connected components of $G-M$,
    arbitrarily pick one redundant $T\in\mathcal{T}$ (which exists as shown in 3.8) and output $(G-V(T),k,M,\mathcal{H})$.
@@
The fact that the rule can be applied in polynomial time directly follows from 3.8. Also, as the graph $G^{\prime}$ obtained after applying the rule is an induced subgraph of $G\in\mathcal{G}$, we have $G^{\prime}\in\mathcal{G}$.
@@
Let $x_{1},x_{2}\in X\setminus S^{\prime}$ be distinct vertices, and suppose by contradiction that there are distinct $T_{1},T_{2}\in\mathcal{T}\setminus\{T\}$ such that $S$ does not intersect $T_{1}$ nor $T_{2}$. There exist $a_{1}\in N_{T_{1}}(x_{1})$ and $b_{1}\in N_{T_{1}}(x_{2})$ as $N_{M}(T_{1})=X$, and because $T_{1}$ is connected there exists a path from $a_{1}$ to $b_{1}$ in $T_{1}$ (observe that we may have $a_{1}=b_{1}$). Similarly, we define $a_{2},b_{2}$ and a path joining them in $T_{2}$. Then we have a cycle not hit by $S^{\prime}$, a contradiction. ∎
So if $|X\setminus S^{\prime}|\geq 2$, there are at least $|X|$ vertices of $S^{\prime}$ in the trees of $\mathcal{T}$. Let denote $Z$ this set. The set $S=S^{\prime}\setminus Z\cup X$ is then a feedback vertex set of $G$ with $|S|\leq|S^{\prime}|$ as wanted.
Second case::
@@
    Otherwise, the graph $G[T\cup\{x\}]$ has no cycle. The cycle $C$ of $G-S^{\prime}$ necessarily uses some vertices of $T$ and we have $N_{M}(T)\setminus S^{\prime}=\{x\}$ and $G[T\cup\{x\}]$ without cycle. This configuration is not possible if $T$ is a connected component of $G-M$, so we are in the case where there exists $r$ a common parent for the roots of the trees in $\mathcal{T}$. The cycle $C$ necessarily contains a path between $x$ and $r$ with inner vertices in $T$. Note that all the $|X|+1\geq 2$ trees $T^{\prime}\in\mathcal{T}\setminus\{T\}$ allow such path between $x$ and $r$. So $S^{\prime}$ contains at least one vertex in some $T^{\prime}\in\mathcal{T}\setminus\{T\}$ and by taking $v$ one such vertex we can take $S=(S^{\prime}\setminus\{v\})\cup\{x\}$ and reach the same conclusion as above.
Third case::
in the remaining case where $X\subseteq S^{\prime}$ we can take $S=S^{\prime}$.
Hence the input and output instances are indeed equivalent. ∎
Definition 3.11.
Given an instance $(G,k,M,\mathcal{H})$, we say that (KR4) applies on $(G,k,M,\mathcal{H})$ if there exists $X$ and $\mathcal{T}$ such that Rule (KR4) can be applied on $(G,k,M,\mathcal{H})$, $X$ and $\mathcal{T}$.
Observe that in (KR4), we consider that the sets $X$ and $\mathcal{T}$ are given, whereas in Algorithm 1 line 2, we have to find these sets. Thus, we need the following lemma.
@@
Given an instance $(G,k,M,\mathcal{H})$, deciding if Rule (KR4) can applies on $(G,k,M,\mathcal{H})$ (and finding $X$ and $\mathcal{T}$ if it is the case) can be done in polynomial time.
@@
We first compute for each connected component $T$ of $G-M$ their neighborhood in $M$ (i.e., $N_{M}(T)$). If there is a set $X$ such that the set $\mathcal{T}_{X}$ of connected components $T$ in $G_{M}$ such that $N_{M}(T)=X$ is large enough, that is $|\mathcal{T}_{X}|\geq|X|+2$, we are done. Finding a family $\mathcal{T}$ for the second version of the rule, where we do not consider connected components of $G-M$ anymore but trees under a common parent $r$, can be done in a similar way that the first case, testing any $r\in V(G-M)$ as the potential common parent. These operations can be performed in polynomial time. ∎
@@ 3.4. Properties of the kernelized instances
The goal of this section is to prove that for an instance $(G,k,M,\mathcal{H})$ for which the kernelization rules do not apply anymore (meaning that we reach Line 5 in Algorithm 1), the size of a subtree $T$ of $G-M$ is strongly related to $\operatorname{\mu}(T)$. (Recall that $\operatorname{\mu}(T)=\max(d_{M}(T),b_{\overline{M}}(T))$.)
@@
For every $K_{r,r}$-free $G\in\mathcal{G}$, every set $A\subseteq V(G)$, and every family $\mathcal{T}$ of disjoint non-adjacent subtrees of $G-A$, and every $x\in\mathbb{R}$, if $|\mathcal{T}|\geq xf_{1}(r)|A|^{\alpha}$ then there exists $X\subseteq A$ such that at least $x$ subtrees $T\in\mathcal{T}$ satisfy $N_{A}(T)=X$. Moreover suppose that for every $T\in\mathcal{T}$ we have $d_{A}(T)\leq p$ and $|T|\leq m$, then if $|\mathcal{T}|\geq xf_{2}(r,p,m)|A|$, there exists $X\subseteq A$ such that at least $x$ subtrees $T\in\mathcal{T}$ satisfy $N_{A}(T)=X$.
@@
The results are obtained from the definition of 1.4 and using the pigeonhole principle. ∎
Recall that given a subtree $T$ of $G-M$, $\partial_{\overline{M}}(T)$ denotes the set $\{v\in V(T),~{}N(v)\not\subseteq M\cup V(T)\}$, and $b_{\overline{M}}(T)$ denotes the size of this set.
We are now ready to bound the degree of the subtrees of $G-M$:
@@
Consider an instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS such that neither the rule (KR1) nor (KR4) applies. For any subtree $T$ of $G-M$ and any vertex $v$ of $T$, we have
@@
Remember that we chose an arbitrary root for $T$, and $T_{v}$ denote the subtree of $T$ with root $v$. For bounding the number of neighbors of $v$ in $T$, it suffice to bound the number of children of $v$. For this, we first use that, by definition, the number of children $u$ of $v$ such that $V(T_{u})\cap\partial_{\overline{M}}(T)\neq\emptyset$ is bounded by $b_{\overline{M}}(T)$. Let $N$ be the set of remaining children, meaning the set of children $u$ such that $T_{u}\cap\partial_{\overline{M}}(T)=\emptyset$. Applying 3.13, with the family $(T_{u})_{u\in N}$, the vertex set $N_{M}(T)$ and $x=d_{M}(T)+2$, if $|N|>(d_{M}(T)+2)f_{1}(r)d_{M}(T)^{\alpha}$ then there would be a set $X\subseteq N_{M}(T)$ such that $d_{M}(T)+2$ vertices $u\in N$ satisfy $d_{M}(T_{u})=X$. Moreover $X\neq\emptyset$ as otherwise the rule (KR1) would have applied. But then Rule (KR4) would apply. Hence $|N|=\mathcal{O}(f_{1}(r)d_{M}(T)^{\alpha+1})$. ∎
@@
A subtree $T$ of $G-M$ is weakly connected to $M$ if $G[T\cup\{u\}]$ is acyclic for every $u$ in $M$ (i.e. $d_{T}(u)\leq 1$ for all $u\in M$). A subtree $T$ of $G-M$, rooted at a vertex $v$, is sharp w.r.t. $M$ if $T$ is not weakly connected to $M$ but for every children $u$ of $v$, $T_{u}$ is weakly connected to $M$.
@@
Consider an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1) and (KR2) applies. For any subtree $T$ in $G-M$ that is weakly connected to $M$, we have that $|T|\leq\operatorname{\mu}(T)$.
@@
  • $d_{G}(v)=2$.
Observe that the leaves of $T$ are either of the first type or the second. Let $l$ be the number of leaves of $T$, there are $b_{\overline{M}}(T)$ vertices of the first type, and at most $d_{M}(T)$ vertices of the second type, so $l\leq b_{\overline{M}}(T)+d_{M}(T)$. Moreover the number of vertices of the third type is at most $l$. Let denote $T_{4}$ the vertices of the fourth type. Observe that the connected components of $G[T_{4}]$ are paths of size at most $3$ (as otherwise (KR2) would applies). Replacing such a connected component by an edge between the neighbors of the endpoints of the path would result in a tree with vertices $T\setminus T_{4}$ and whose number of edges (which is at most $|T\setminus T_{4}|-1)$ is an upper bound on the number of connected components of $G[T_{4}]$. So $|T_{4}|\leq 3|T\setminus T_{4}|\leq 6l$, and finally $T$ has less than $8l=8d_{M}(T)+8b_{\overline{M}}(T)$ vertices in total. ∎
We now consider sharp subtrees. Notice that in such a tree, $v$ has bounded degree (by 3.14), and the subtrees below $v$ have bounded size (by 3.16). This leads to the following corollary.
@@
Consider an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1), (KR2) and (KR4) applies. Given a sharp subtree $T$ of $G-M$, we have $|T|=O\left(f_{1}(r)d_{M}(T)^{\alpha}\operatorname{\mu}(T)^{2}\right)$.
@@
3.5. Kernelizing when a big tree is found.
When reaching line 5 of Algorithm 1, we call the method Partitionning-Algorithm  which tries to build a special partition of $G-M$. As we will see later, one output of this procedure is a failure (called case 1) where a “big” (whose size is too large with respect to $d_{M}(T)$ and $b_{\overline{M}}(T)$) tree $T$ is found in $G-M$. In this section, we explain how we can get rid of such a big tree.
@@
    Consider an instance $(G,k,M,\mathcal{H})$, with a subtree $T$ of $G-M$ which contains $d_{M}(T)+b_{\overline{M}}(T)$ vertex disjoint paths of length at least $1$ and whose endpoints are all adjacent to some vertex $u\in N_{M}(T)$. Then output $\left(G-u,k-1,M\setminus\{u\},\mathcal{H}-\{u\}\right)$ if $k\geq 1$ (or a trivial no instance otherwise).
@@
Again, as the graph $G^{\prime}$ obtained after applying the rule is an induced subgraph of $G$, the rule preserves the property of being a $K_{r,r}$-free $s$-string graph. For the safeness it is sufficient to prove that if $G$ has a feedback vertex set of size at most $k$, then it has a solution $S^{\prime}$ of size at most $k$ containing $u$. Let $S$ be a feedback vertex set of size at most $k$, and suppose it does not contains $u$. Then $S$ has at least a vertex in each of the $d_{M}(T)+b_{\overline{M}}(T)$ considered disjoints paths in $T$. So the set $S^{\prime}=(S\setminus T)\cup N_{M}(T)\cup\partial_{\overline{M}}(T)$ satisfies $|S^{\prime}|\leq|S|\leq k$ and contains $u$. Moreover this set is a feedback vertex set of $G$ : a cycle in $G-S^{\prime}$ would contains vertices of $T$ as $S$ is a feedback vertex set of $G$, but then, as $T$ is a tree, such a cycle would need to intersect the set $N_{G}(T)=N_{M}(T)\cup\partial_{\overline{M}}(T)\subseteq S^{\prime}$, which is a contradiction. ∎
@@
There is a multivariate polynomial $p_{1}$ with $p_{1}(x,y)=\mathcal{O}(f_{1}(x)y^{6+\alpha})$ such that for every instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS where none of the rules (KR1), (KR2), and (KR4) applies, and for every subtree $T$ of $G-M$ such that the Rule (KR5) does not apply, we have $|T|\leq p_{1}(r,\operatorname{\mu}(T))$.

### v2

@@ 3.3. Kernelization rules
Here we provide the four kernelization rules (KR1), (KR2), (KR3), and (KR4) that Algorithm 1 tries to apply at line 2. Each of these rules takes as input an instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS and outputs a single instance $(G^{\prime},k^{\prime},M^{\prime},\mathcal{H}^{\prime})$. Such a rule is said to be safe if:
  • For an instance of $(r,\mathcal{G})$-Ann-FVS the rule returns an instance of $(r,\mathcal{G})$-Ann-FVS (in particular the graph of the output instance is a $K_{r,r}$-free graph in $\mathcal{G}$), and
@@
Given any instance $(G,k,M,\mathcal{H})$, recall that every connected component of GMG-M is a tree. We root each of them at an arbitrary vertex. We define a subforest of GMG-M as a subset $T\subseteq V(G-M)$ and say the set is a subtree of GMG-M if $G[T]$ is a tree. Given a vertex $v$ of a connected component $T$ in GMG-M, we define the subtree $T_{v}$ of GMG-M as the connected component of $v$ in GMuG-M-u, where $u$ is the parent of $v$, if any. If $v$ is the root of $T$, then $T_{v}=T$. In any case $T_{v}$ is rooted at $v$. Given $X\subseteq V(G)$, we denote $\mathcal{H}-X=\{H\in\mathcal{H}\mid H\cap X=\emptyset\}$. Given a subtree $T$ of GMG-M, $\partial_{\overline{M}}(T)$ denotes the set $\{v\in T,~N(v)\not\subseteq M\cup T\}$, and $b_{\overline{M}}(T)$ denotes the size of this set. We also denote $\operatorname{\mu}(T)=\max(d_{M}(T),b_{\overline{M}}(T))$.
We start with two basic reduction rules often used to deal with FVS and that allow to get rid of vertices of degree 1 and arbitrarily long paths of vertices of degree 2. Notice that we only apply here the reduction to vertices in GMG-M.
@@
    Given an instance $(G,k,M,\mathcal{H})$, if there exists a vertex $v\in V(G-M)$ of degree $d(v)\leq 1$, output $(G-\{v\},k,M,\mathcal{H})$.
@@
    Given an instance $(G,k,M,\mathcal{H})$, if there exists a path quvwquvw in V(GM)V(G-M) such that the four vertices have degree $2$ in $G$ (hence $d_{M}(u)=d_{M}(v)=0$), output $(G^{\prime},k,M,\mathcal{H})$, where $G^{\prime}$ is the graph obtained from $G$ by contracting the edge $uv$.
@@
Notice that the vertices considered in the two rules belong to GMG-M, in particular they do not belong to any member of $\mathcal{H}$ (which are subsets of $M$). Contracting an edge with endpoints of degree two preserves being $K_{r,r}$-free and does not modify the size of the minimum feedback vertex sets. Moreover the obtained graph is still in $\mathcal{G}$ by definition of a nice class. The output is then an instance of $(r,\mathcal{G})$-Ann-FVS equivalent to the input. The running time claim is immediate.
@@
Observe that we did not use the condition in (KR2) that the endpoints of the considered path have degree at most $2$ in $G$. However it will later be used in 3.32 to bound the size of $M$ during the execution of the algorithm.
The next rule ensures that the vertices outside $M$ have a small neighborhood in $M$. It increases the size of $M$, but in a controlled manner as we will see later in 3.32.
@@
    Given an instance $(G,k,M,\mathcal{H})$, if there is a vertex $v\in V(G-M)$ such that $d_{M}(v)\geq t=2d(r)$, with $d(r)$ the value defined in 1.5, output $(G,k,M\cup\{v\},\mathcal{H})$.
@@
The fourth kernelization rule deletes unnecessary trees of GMG-M. Consider an instance $(G,k,M,\mathcal{H})$ where none of the previous rules applies. Given a family $\mathcal{T}$ of disjoint subtrees of GMG-M, a tree $T\in\mathcal{T}$ is redundant (for $\mathcal{T}$) if for all $v\in M$ such that $d_{T}(v)\geq 2$, there exists $T^{\prime}\in\mathcal{T}$ with $T^{\prime}\neq T$ such that $d_{T^{\prime}}(v)\geq 2$.
@@
Consider a set $X\subseteq M$ and a set $\mathcal{T}$ of more than $|X|$ subtrees of GMG-M such that for every $T\in\mathcal{T}$, $N_{M}(T)=X$. Then, there exists a redundant tree for $\mathcal{T}$, and it can be found in polynomial time.
@@
For any $x\in X$, if there is at least one subtree $T\in\mathcal{T}$ such that $d_{T}(x)\geq 2$, we arbitrarily pick one of them and call it $T^{(x)}$. Now, let $T^{\prime}$ be one of the remaining trees in $\mathcal{T}$. Such a tree exists as $|\mathcal{T}|\geq|X|+1$, and there are at most $|X|$ trees of the form $T^{(x)}$, for some $x\in X$. It is immediate from the definition that $T^{\prime}$ is redundant. ∎
Recall that we consider a fixed child-parent orientation in the forest GMG-M. In what follows, we say that a subset $F\subseteq V(G-M)$ is a downward-closed subtree (or subforest when the set is not necessarily connected) of GMG-M when for any $v\in F$ and any children $u$ of $v$, $u\in F$. The fourth kernelization rule is as follows:
@@
    Given an instance $(G,k,M,\mathcal{H})$, a set $X\subseteq M$ with $|X|\geq 1$ and a set $\mathcal{T}$ of at least $|X|+2$ disjoint downward-closed subtrees of GMG-M such that:
@@
    • either all the roots of the trees in $\mathcal{T}$ have a common parent $r$, or $\mathcal{T}$ consists only in connected components of GMG-M,
    arbitrarily pick one redundant $T\in\mathcal{T}$ (which exists as shown in 3.7) and output $(G-V(T),k,M,\mathcal{H})$.
@@
The fact that the rule can be applied in polynomial time directly follows from 3.7. Also, as the graph $G^{\prime}$ obtained after applying the rule is an induced subgraph of $G\in\mathcal{G}$, we have $G^{\prime}\in\mathcal{G}$.
@@
[…]
So if $|X\setminus S^{\prime}|\geq 2$, there are at least $|X|$ vertices of $S^{\prime}$ in the trees of $\mathcal{T}$. Let us denote $Z$ this set. The set $S=S^{\prime}\setminus Z\cup X$ is then a feedback vertex set of $G$ with $|S|\leq|S^{\prime}|$ as wanted.
Second case:
@@
    Otherwise, the graph $G[T\cup\{x\}]$ has no cycle. The cycle $C$ of $G-S^{\prime}$ necessarily uses some vertices of $T$ and we have $N_{M}(T)\setminus S^{\prime}=\{x\}$ and $G[T\cup\{x\}]$ without cycle. This configuration is not possible if $T$ is a connected component of GMG-M, so we are in the case where there exists $r$ a common parent for the roots of the trees in $\mathcal{T}$. The cycle $C$ necessarily contains a path between $x$ and $r$ with inner vertices in $T$. Note that all the $|X|+1\geq 2$ trees $T^{\prime}\in\mathcal{T}\setminus\{T\}$ allow such path between $x$ and $r$. So $S^{\prime}$ contains at least one vertex in some $T^{\prime}\in\mathcal{T}\setminus\{T\}$ and by taking one such vertex $v$ we can set $S=(S^{\prime}\setminus\{v\})\cup\{x\}$ and reach the same conclusion as above.
Third case:
in the remaining case, where $X\subseteq S^{\prime}$, we can take $S=S^{\prime}$.
Hence, the input and output instances are indeed equivalent. ∎
Observe that Algorithm 1 requires a routine for checking if rule (KR4) applies, before applying it (with 3.8).
@@
Given an instance $(G,k,M,\mathcal{H})$, deciding if Rule (KR4) can be applied, and finding $X$ and $\mathcal{T}$ if it is the case, can be done in polynomial time.
@@
We first compute for each connected component $T$ of GMG-M their neighborhood in $M$ (i.e., $N_{M}(T)$). If there is a set $X$ such that the set $\mathcal{T}_{X}$ of connected components $T$ in $G_{M}$ such that $N_{M}(T)=X$ is large enough, that is $|\mathcal{T}_{X}|\geq|X|+2$, we are done. Finding a family $\mathcal{T}$ for the second version of the rule, where we do not consider connected components of GMG-M anymore but trees under a common parent $r$, can be done in a similar way by first testing any $r\in V(G-M)$ as the potential common parent. These operations can be performed in polynomial time. ∎
@@ 3.4. Properties of the kernelized instances
The goal of this section is to prove that for an instance $(G,k,M,\mathcal{H})$ for which the kernelization rules do not apply anymore (meaning that we reach line 5 in Algorithm 1), the size of a subtree $T$ of GMG-M is strongly related to $\operatorname{\mu}(T)$. (Recall that $\operatorname{\mu}(T)=\max(d_{M}(T),b_{\overline{M}}(T))$.)
@@
For every $K_{r,r}$-free $G\in\mathcal{G}$, every set $A\subseteq V(G)$, and every family $\mathcal{T}$ of disjoint non-adjacent subtrees of GAG-A, and every integer $x$, if $|\mathcal{T}|\geq xf_{1}(r)|A|^{\alpha}$ then there exists $X\subseteq A$ such that at least $x$ subtrees $T\in\mathcal{T}$ satisfy $N_{A}(T)=X$. Moreover, suppose that for every $T\in\mathcal{T}$ we have $d_{A}(T)\leq p$ and $|T|\leq m$, then if $|\mathcal{T}|\geq xf_{2}(r,p,m)|A|$, there exists $X\subseteq A$ such that at least $x$ subtrees $T\in\mathcal{T}$ satisfy $N_{A}(T)=X$.
@@
The results are obtained from 1.4 by using the pigeonhole principle. ∎
Recall that given a subtree $T$ of GMG-M, $\partial_{\overline{M}}(T)$ denotes the set $\{v\in V(T),~N(v)\not\subseteq M\cup V(T)\}$, and $b_{\overline{M}}(T)$ denotes the size of this set.
We are now ready to bound the degree of the subtrees of GMG-M:
@@
Consider an instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS such that neither Rule (KR1) nor (KR4) applies. For any subtree $T$ of GMG-M and any vertex $v$ of $T$, we have
@@
Remember that we chose an arbitrary root for $T$, and $T_{v}$ denotes the subtree of $T$ with root $v$. For bounding the number of neighbors of $v$ in $T$, it suffice to bound the number of children of $v$. For this, we first use that the number of children $u$ of $v$ such that $V(T_{u})\cap\partial_{\overline{M}}(T)\neq\emptyset$ is bounded by $b_{\overline{M}}(T)$. Let $N$ be the set of remaining children, meaning the set of children $u$ such that $T_{u}\cap\partial_{\overline{M}}(T)=\emptyset$. Applying 3.11, with the family $(T_{u})_{u\in N}$, the vertex set $N_{M}(T)$ and $x=d_{M}(T)+2$, if $|N|>(d_{M}(T)+2)f_{1}(r)d_{M}(T)^{\alpha}$ then there would be a set $X\subseteq N_{M}(T)$ such that $d_{M}(T)+2$ vertices $u\in N$ satisfy $d_{M}(T_{u})=X$. Moreover $X\neq\emptyset$ as otherwise Rule (KR1) would have applied. But then Rule (KR4) would apply. Hence $|N|=\mathcal{O}(f_{1}(r)d_{M}(T)^{\alpha+1})$. ∎
@@
A subtree $T$ of GMG-M is weakly connected to $M$ if $G[T\cup\{u\}]$ is acyclic for every $u$ in $M$ (i.e. $d_{T}(u)\leq 1$ for all $u\in M$). A subtree $T$ of GMG-M, rooted at a vertex $v$, is sharp w.r.t. $M$ if $T$ is not weakly connected to $M$ but for every child $u$ of $v$, $T_{u}$ is weakly connected to $M$.
@@
Consider an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1) and (KR2) applies. For any subtree $T$ in GMG-M that is weakly connected to $M$, we have that $|T|\leq 16\operatorname{\mu}(T)$.
@@
  4. (4)
    $d_{T}(v)=2$.
Observe that the leaves of $T$ are either of the first type or the second. Let $l$ be the number of leaves of $T$, there are $b_{\overline{M}}(T)$ vertices of the first type, and at most $d_{M}(T)$ vertices of the second type, so $l\leq b_{\overline{M}}(T)+d_{M}(T)$. Moreover the number of vertices of the third type is at most $l$. Let denote by $Z$ the vertices of the fourth type. Observe that the connected components of $G[Z]$ are paths of size at most $3$ (as otherwise (KR2) would apply). Replacing such a connected component by an edge between the neighbors of the endpoints of the path would result in a tree with vertices $T\setminus Z$ and whose number of edges (which is at most $|T\setminus Z|-1)$ is an upper bound on the number of connected components of $G[Z]$. So $|Z|\leq 3|T\setminus Z|\leq 6l$, and finally $T$ has less than $8l=8d_{M}(T)+8b_{\overline{M}}(T)\leq 16\operatorname{\mu}(T)$ vertices in total. ∎
We now consider sharp subtrees. Notice that in such a tree, $v$ has bounded degree (by 3.12), and the subtrees below $v$ have bounded size (by 3.14). This leads to the following corollary.
@@
Consider an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1), (KR2) and (KR4) applies. Given a sharp subtree $T$ of GMG-M, we have $|T|=O\left(f_{1}(r)d_{M}(T)^{\alpha}\operatorname{\mu}(T)^{2}\right)$.
@@
3.5. Kernelization when a big tree is found.
When reaching line 5 of Algorithm 1, we call the routine Partitionning-Algorithm  which tries to build a special partition of GMG-M. As we will see later, one output of this procedure is a failure (called Case 1) where a “big” (whose size is too large with respect to $d_{M}(T)$ and $b_{\overline{M}}(T)$) tree $T$ is found in GMG-M. In this section, we explain how we can get rid of such a big tree.
@@
    Consider an instance $(G,k,M,\mathcal{H})$, with a subtree $T$ of GMG-M which contains $d_{M}(T)+b_{\overline{M}}(T)$ vertex disjoint paths of length at least $1$ and whose endpoints are all adjacent to some vertex $u\in M$. Then output $\left(G-u,k-1,M\setminus\{u\},\mathcal{H}-\{u\}\right)$ if $k\geq 1$ and a trivial No-instance otherwise. [^8] Recall that for a set $\subseteq V(G), V\mathcal{H}-V$ is defined (at the beginning of Subsection 3.3) as the set of those members of $\mathcal{H}$ not intersected by $V$, i.e., $\{H\in\mathcal{H}\mid H\cap V=\emptyset\}$.
@@
Again, as the graph $G^{\prime}$ obtained after applying the rule is an induced subgraph of $G$, the rule preserves the property of being a $K_{r,r}$-free graph in the considered graph class. For the safeness it is sufficient to prove that if $G$ has a minimum feedback vertex set of size at most $k$, then it has a solution $S^{\prime}$ of size at most $k$ containing $u$. Let $S$ be a minimum feedback vertex set, and suppose it does not contain $u$. Then $S$ has at least a vertex in each of the $d_{M}(T)+b_{\overline{M}}(T)$ considered disjoints paths in $T$. So the set $S^{\prime}=(S\setminus T)\cup N_{M}(T)\cup\partial_{\overline{M}}(T)$ satisfies $|S^{\prime}|\leq|S|\leq k$ and contains $u$. Moreover this set is a feedback vertex set of $G$ : a cycle in $G-S^{\prime}$ would contain vertices of $T$ as $S$ is a feedback vertex set of $G$, but then, as $T$ is a tree, such a cycle would need to intersect the set $N_{G}(T)=N_{M}(T)\cup\partial_{\overline{M}}(T)\subseteq S^{\prime}$, which is a contradiction. ∎
@@
There is a multivariate polynomial $p_{1}$ with $p_{1}(x,y)=\mathcal{O}(f_{1}(x)y^{6+\alpha})$ such that for every instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS where none of the rules (KR1), (KR2), and (KR4) applies, and for every subtree $T$ of GMG-M such that the rule (KR5) does not apply, we have $|T|\leq p_{1}(r,\operatorname{\mu}(T))$.